# Lossy Compression with Data, Perception, and Classification Constraints

### v2

@@ Appendix E Convexity of RDC function
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\sum_{\alpha=1}^{k}\mathbf{w}[\alpha]\cdot{\mathbf{d}_{\alpha}^{T}\mathbf{p}_{\alpha}},$ (15)
@@
Consider the mapping $\mathbf{p}_{\alpha}\to(\mathbf{p}_{\alpha},H(\mathbf{p}_{\alpha}),H(\mathbf{Tp}_{\alpha}),\mathbf{d}_{\alpha}^{T}\mathbf{p}_{\alpha})$, and let $\mathcal{S}$ be the set of all points of $(\mathbf{p}_{\alpha},H(\mathbf{p}_{\alpha}),H(\mathbf{Tp}_{\alpha}),\mathbf{d}_{\alpha}^{T}\mathbf{p}_{\alpha})$ for $\mathbf{p}_{\alpha}\in\Delta_{n}$.
Lemma 1.
@@
Proof.
On the one hand, relations (12)-(15) represent $(\mathbf{p},\rho,\eta,\delta)$ as a convex combination with weights $\mathbf{w}[\alpha]$ of the $k$ points $(\mathbf{p}_{\alpha},H(\mathbf{p}_{\alpha}),H(\mathbf{Tp}_{\alpha}),\mathbf{d}_{\alpha}^{T}\mathbf{p}_{\alpha})$ all belonging to $\mathcal{S}$. On the other hand, every point of the convex hull $\mathcal{C}$ of $\mathcal{S}$ is a convex combination of a finite number of points of $\mathcal{S}$, hence it is of the form $(\mathbf{p},\rho,\eta,\delta)$. ∎
@@
$(D_{\lambda},C_{\lambda})=\lambda(D_{1},C_{1})+(1-\lambda)(D_{2},C_{2}),$
$\rho_{\lambda}=\lambda\rho_{1}+(1-\lambda)\rho_{2}.$
@@
$F(D_{\lambda},C_{\lambda})$ $=\sup\{\rho|(\mathbf{q},\rho,C_{\lambda},D_{\lambda})\in\mathcal{C}\}$
$\geq\rho_{\lambda}$
$=\lambda F(D_{1},C_{1})+(1-\lambda)F(D_{2},C_{2}),$
@@
$R(\lambda D_{1}$ $+(1-\lambda)D2,\lambda C_{1}+(1-\lambda)C_{2})$
$\leq\lambda R(D_{1},C_{1})+(1-\lambda)R(D_{2},C_{2}),$
@@ F-A Optimality with strong asymptotical constraints
Definition 1 ((Information) rate function [15]).
@@
Definition 2 (Asymptotically achievable with common randomness [15]).
@@ F-B Achievability with Weak Asymptotical Constraints and Optimality of the RDC Function
Definition 3 (Weak asymptotically achievable with common randomness).

### v3

@@ Appendix E Convexity of RDC function
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\sum_{\alpha=1}^{k}\mathbf{w}[\alpha]\cdot{\mathbf{d}_{\alpha}^{\top}\mathbf{p}_{\alpha}},$ (15)
@@
Consider the mapping $\mathbf{p}_{\alpha}\to(\mathbf{p}_{\alpha},H(\mathbf{p}_{\alpha}),H(\mathbf{Tp}_{\alpha}),\mathbf{d}_{\alpha}^{\top}\mathbf{p}_{\alpha})$, and let $\mathcal{S}$ be the set of all points of $(\mathbf{p}_{\alpha},H(\mathbf{p}_{\alpha}),H(\mathbf{Tp}_{\alpha}),\mathbf{d}_{\alpha}^{\top}\mathbf{p}_{\alpha})$ for $\mathbf{p}_{\alpha}\in\Delta_{n}$.
Lemma 1
@@
Proof:
On the one hand, relations (12)-(15) represent $(\mathbf{p},\rho,\eta,\delta)$ as a convex combination with weights $\mathbf{w}[\alpha]$ of the $k$ points $(\mathbf{p}_{\alpha},H(\mathbf{p}_{\alpha}),H(\mathbf{Tp}_{\alpha}),\mathbf{d}_{\alpha}^{\top}\mathbf{p}_{\alpha})$ all belonging to $\mathcal{S}$. On the other hand, every point of the convex hull $\mathcal{C}$ of $\mathcal{S}$ is a convex combination of a finite number of points of $\mathcal{S}$, hence it is of the form $(\mathbf{p},\rho,\eta,\delta)$. ∎
@@
$(D_{\lambda},C_{\lambda})$ $\triangleq\lambda(D_{1},C_{1})+(1-\lambda)(D_{2},C_{2}),$
$\rho_{\lambda}$ $\triangleq\lambda\rho_{1}+(1-\lambda)\rho_{2}.$
@@
$F(D_{\lambda},C_{\lambda})$ $=\sup\{\rho|(\mathbf{q},\rho,C_{\lambda},D_{\lambda})\in\mathcal{C}\}\geq\rho_{\lambda}=\lambda F(D_{1},C_{1})+(1-\lambda)F(D_{2},C_{2}),$
@@
$R(\lambda D_{1}$ $+(1-\lambda)D2,\lambda C_{1}+(1-\lambda)C_{2})\leq\lambda R(D_{1},C_{1})+(1-\lambda)R(D_{2},C_{2}),$
@@ F-A Optimality with strong asymptotical constraints
Definition 1 ((Information) rate function [16])
@@
Definition 2 (Asymptotically achievable with common randomness [16])
@@ F-B Achievability with Weak Asymptotical Constraints and Optimality of the RDC Function
Definition 3 (Weak asymptotically achievable with common randomness)